OpenStudy (anonymous):

prove that coss^2A-sinn^A = 2cos^2A-1=1-2sin^2A

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

I know that cos^2A- 1- sin^2A

OpenStudy (anonymous):

=*

OpenStudy (aravindg):

easy to get both results substitute sin^2x and cos^x from sin^ x+cos^2 x=1

OpenStudy (anonymous):

we need to get rid of sin for the first bit I think

OpenStudy (aravindg):

yep ! for example for first one cos^2 a-sin^2 a=cos^2 a -(1-sin^2 a )=cos^2 a +cos^2 a -1 =2cos^2 a-1

OpenStudy (anonymous):

ahI got the first part

OpenStudy (anonymous):

yay

OpenStudy (aravindg):

now can you try second part? :)

OpenStudy (anonymous):

yes thanks

OpenStudy (aravindg):

yw :)